# Nonreciprocal entanglement in cavity magnomechanics exploiting chiral cavity-magnon coupling

### v3

@@ II The Model
We start with the case where the CW mode is driven. The CCW mode is essentially decoupled from the tripartite system of the CW, magnon, and mechanical modes, because of $g_{\circlearrowright}>0$ and $g_{\circlearrowleft}=0$. The system is then reduced to the tripartite CMM system studied in Li2018 , typically for a three-dimensional microwave cavity Zuo2023 . Following Li2018 , we adopt a red-detuned microwave drive field with respect to the magnon mode with detuning $\tilde{\Delta}_{m}\approx\omega_{b}$ (Fig. 1(b)), which is responsible for significantly cooling the low-frequency mechanical mode to eliminate the dominant thermal noise entering the system. The cooling interaction optimally works in the resolved-sideband limit of $\kappa_{m}\ll\omega_{b}$, which is well satisfied for the CMM experiments using YIG Zhang2016 ; Potts2021 ; Shen2022 ; Shen23 . A precondition for seeing entanglement in the system is that the mechanical mode should be cooled into its quantum ground state Li2018 ; Li2019A ; Li2020 . However, with only the effective magnomechanical cooling (beam-splitter) interaction, it is insufficient to produce entanglement. The magnomechanical parametric down conversion (PDC) interaction should be activated. Typically, this is realized by driving the magnon mode with a blue-detuned microwave field, which, however, can easily cause the system to be unstable. We thus adopt a sufficiently strong red-detuned drive field Vitali2007 , such that the weak-coupling condition $G_{m}\ll\omega_{b}$ for taking the rotating-wave (RW) approximation to obtain the magnomechanical beam-splitter interaction $\propto m^{\dagger}b+mb^{\dagger}$ is broken, where $b$ denotes the mechanical mode and $b=(q+ip)/\!\sqrt{2}$. Consequently, the counter-RW terms, which correspond to the PDC interaction $\propto m^{\dagger}b^{\dagger}+mb$, start to play the role in creating magnomechanical entanglement. The entanglement is further distributed to the circulating cavity mode when the cavity is resonant with the Stokes mechanical sideband, i.e., $\Delta_{a}\approx-\omega_{b}$, yielding the entanglement between the CW microwave mode with the magnon (mechanical) mode, as shown in Fig. 2(a) (Fig. 2(b)). The entanglements are maximized around the optimal detunings as depicted in Fig. 1(b), i.e., the CW cavity and magnon modes are respectively resonant with the Stokes and anti-Stokes sidebands of the drive field.
@@
We have employed feasible parameters Zhang2016 ; Li2018 ; Potts2021 ; Shen2022 ; Shen23 ; Bourhill2023 : $\omega_{a}/2\pi\approx\omega_{m}/2\pi\approx\omega_{0}/2\pi=10$ GHz, $\omega_{b}/2\pi=10$ MHz, $\kappa_{m}/2\pi=1$ MHz, $\gamma_{b}/2\pi=10^{2}$ Hz, $\kappa_{a,i}/2\pi=0.2$ MHz, $\kappa_{a,e}/2\pi=2.8$ MHz ($4.8$ MHz), $g_{\circlearrowright}/2\pi=4$ MHz ($8$ MHz), and $T=10$ mK in getting Fig. 2(a) (Fig. 2(b)). These parameters lead to a complex steady-state average of $\left\langle m\right\rangle$, cf. Eq. (5). Therefore, Fig. 2(a) (Fig. 2(b)) is plotted with a fixed amplitude of the effective coupling $|G_{m}|/2\pi=4.0$ MHz ($2.5$ MHz). Figure 2(c) shows the drive power required to generate considerable microwave-phonon entanglement for different values of the cavity decay rate. Clearly, one can increase the drive power to compensate the cavity losses in such a torus-shaped microwave cavity. Figure 2(d) plots the entanglements $E_{a_{\circlearrowright}m}$ and $E_{a_{\circlearrowright}b}$ versus the mechanical damping rate $\gamma_{b}$. Both the entanglements can survive even for a very large damping rate $\gamma_{b}/2\pi\sim 10^{4}$ Hz, corresponding to the mechanical quality factor $Q_{b}\sim 10^{3}$, which can be easily achieved in the CMM experiments Zuo2023 . Note that the results shown in Fig. 2 are in the steady state, which is guaranteed by the negative eigenvalues (real parts) of the drift matrix $A$.
@@
For the case where the CCW cavity mode is driven, due to the absence of the CCW mode-magnon coupling, i.e., $g_{\circlearrowleft}=0$, the magnon mode is effectively not driven, and hence the dispersive magnomechanical coupling is too weak to create any entanglement Zuo2023 , i.e., $E_{a_{\circlearrowleft}m}=E_{a_{\circlearrowleft}b}=0$. Therefore, significant nonreciprocal microwave-magnon and -phonon entanglements are achieved by exploiting the chiral cavity-magnon coupling in the system.
IV nonreciprocal entanglements versus experimental imperfections
In an actual experiment, various imperfections may cause appreciable effects on the nonreciprocity of the entanglements achieved in the ideal case (Sec. III). These include the coupling between the two degenerate counter-propagating cavity modes by backscattering from the YIG sphere or the cavity ports Bourhill2023 , which can be modeled as $\hbar J(a_{\circlearrowright}a_{\circlearrowleft}^{\dagger}+a_{\circlearrowright}^{\dagger}a_{\circlearrowleft})$ with $J$ being the coupling strength, and the residual coupling of the Kittel mode to the CCW circulating mode, i.e., $g_{\circlearrowleft}\ll g_{\circlearrowright}$ but $g_{\circlearrowleft}\neq 0$, due to, e.g., the inaccurate position of the YIG sphere in the cavity. The inclusion of these two effects leads to a new Hamiltonian of the system and linearized QLEs for the quantum fluctuations, which are provided in the Appendix for brevity.
@@
Figure 3 manifests the impacts of the above two imperfections on the nonreciprocity of the entanglements. Specifically, in Fig. 3(a) we show the nonreciprocal microwave-magnon entanglement as a function of the mode coupling strength $J$ under $\chi\equiv g_{\circlearrowleft}/g_{\circlearrowright}=0.1$ ($g_{\circlearrowright}$ is fixed as in Fig. 2(a)). The four curves correspond, respectively, to the entanglement between the magnon mode with the CW or the CCW mode when either the CW or the CCW mode is driven. For the two solid curves with the CW mode under drive, the increase of the mode coupling $J$ leads to the distribution of the entanglement from the CW mode to the CCW mode, because they are coupled via the beam-splitter (state-swap) interaction Li2018 ; Li2019A ; Li2020 , as witnessed by the increasing $E_{a_{\circlearrowleft}m}^{(\circlearrowright)}$ (the declining $E_{a_{\circlearrowright}m}^{(\circlearrowright)}$ ) with the coupling $J$ in Fig. 3(a). By contrast, for the two dashed curves with the CCW mode under drive, the CCW mode-magnon coupling $g_{\circlearrowleft}=0.1g_{\circlearrowright}$ is too weak (under the same power) to efficiently drive the magnon mode to obtain a sufficiently strong coupling $G_{m}$ for creating entanglement. Although the mode coupling also leads to the driving of the CW mode, this driving is typically weak and also inefficient to achieve a strong $G_{m}$ for generating considerable entanglement. Consequently, the entanglements $E_{a_{\circlearrowleft}m}^{(\circlearrowleft)}$ and $E_{a_{\circlearrowright}m}^{(\circlearrowleft)}$ are either zero or negligibly small. Similar results are observed for the microwave-phonon entanglement as shown in Fig. 3(b). We therefore conclude that the nonreciprocity of both the microwave-magnon and -phonon entanglements is robust against the backscattering induced mode coupling, even when the chiral coupling condition $g_{\circlearrowright}\gg g_{\circlearrowleft}=0$ is slightly broken.
@@
In Fig. 4, we study the effect of thermal noises on the nonreciprocal entanglements. In general, the lower the bath temperature is, the stronger the entanglements are. Nevertheless, too low temperature also leads to the emergence of the other undesired entanglements. One thus can raise a bit the temperature, e.g., around 80 mK, to completely kill the unwanted entanglements, at the price of slightly reducing the desired entanglement. The nonreciprocal cavity-magnon and -phonon entanglements can be achieved for the temperature up to 125 mK and 163 mK, respectively, as shown in Figs. 4(a) and 4(b).
Figure 4: Stationary entanglement $E_{a_{\circlearrowright}\nu}^{(\circlearrowright)}$ (thick solid), $E_{a_{\circlearrowleft}\nu}^{(\circlearrowright)}$ (thin solid), $E_{a_{\circlearrowright}\nu}^{(\circlearrowleft)}$ (thick dashed) and $E_{a_{\circlearrowleft}\nu}^{(\circlearrowleft)}$ (thin dashed) versus bath temperature $T$. Here, $\nu=m$ in (a) for microwave-magnon entanglement; $\nu=b$ in (b) for microwave-phonon entanglement. We take $\chi=0.1$ and the other parameters of (a) and (b) are the same as in Figs. 3(c) and 3(d), respectively.
@@
Apart from the nonreciprocal bipartite entanglements studied in the preceding sections, we confirm that the nonreciprocal genuine photon-magnon-phonon tripartite entanglement can also be achieved in the system. The genuine tripartite entanglement is quantified by the minimum residual contangle $R_{\tau}^{\mathrm{min}}$ Adesso2007 , which can be conveniently computed when the CM is achieved Li2018 . In Fig. 5, we study the nonreciprocal tripartite entanglements versus the mode coupling $J$ (Fig. 5(a)) and the coupling ratio $\chi$ (Fig. 5(b)). The four curves represent the tripartite entanglement of the magnon mode, the mechanical mode, and the CW or the CCW mode for the two situations when the CW and CCW modes are respectively driven. The nonreciprocity of the tripartite entanglement is clearly shown to be robust against the mode coupling $J$ and the residual coupling $g_{\circlearrowleft}$ to the CCW mode, as witnessed by the vanishing residual contangle $R_{\tau}^{\mathrm{min}}\approx 0$ of the three unwanted tripartite entanglements.
V Potential applications
The proposed nonreciprocal entanglements find many potential applications, one of which is to realize the channel multiplexing quantum teleportation, as depicted in Fig. 6. In the conventional continuous-variables (CV) quantum teleportation scheme (Fig. 6(a)) cvtele , the teleportation of a quantum state from the sender onto the quantum memory exploits a single channel. By contrast, the nonreciprocal entanglement allows for a channel multiplexing protocol (Fig. 6(b)), where the chiral cavity-magnon coupling enables the microwave photons injecting from the left (right) side to couple with the magnomechanical system in the right (left) cavity and entangle with the magnons/phonons contained therein, but decouple from the magnomechanical system in the left (right) cavity. This allows for the teleportation of the quantum state from the senders at both sides onto their respective quantum memories still using a single channel, thereby increasing the usage efficiency of the channel compared to the conventional teleportation scheme.
Figure 6: (a) Schematic diagram of the conventional CV quantum teleportation. By exploiting the prepared Einstein-Podolsky-Rosen (EPR) state, the “Bell-state” measurement (realized via the homodyne detection) teleports the quantum state from the sender to the quantum memory cvtele . (b) The channel multiplexing quantum teleportation protocol based on the nonreciprocal entanglement proposed in this work. See text for description.
We now discuss the performance of such a channel multiplexing protocol based on, e.g., the generated nonreciprocal microwave-magnon entanglement, and calculate the corresponding teleportation state fidelity. In the CV quantum teleportation, the fidelity of teleporting an input single-mode Gaussian state by exploiting a two-mode Gaussian entangled state is given by Fiurasek2002
@@ V Potential applications
where $V=2V_{in}+\sigma_{z}V_{e}\sigma_{z}^{T}+\sigma_{z}V_{ef}+V_{ef}^{T}\sigma_{z}^{T}+V_{f}$, with $\sigma_{z}=\mathrm{diag}[1,-1]$ being the Pauli-$z$ matrix, $V_{in}$ being the CM of the input state, and $V_{e}$, $V_{f}$, $V_{ef}$ and $V_{fe}$ being $2\times 2$ blocks of the CM of the two-mode entangled state. To be specific, we consider a microwave coherent state as the input state with its CM $V_{in}=\mathrm{diag}[1/2,1/2]$, and the generated microwave-magnon entangled state as the two-mode entangled state. It should be noted that the above microwave-magnon entanglement is between the microwave cavity mode and the magnon mode, while the subsequent “Bell-state” measurement is applied onto the cavity output field and the input coherent state. To estimate the fidelity using Eq. (8), one has to properly define the cavity output mode to have a physically valid CM. The cavity output mode can be defined with a filter function with a properly chosen central frequency $\Omega$ and bandwidth $1/\tau$ (F1 and F2 in Fig. 6(b) denote two filters), given by Genes2008
@@
Under the parameters of Fig. 2(d) and taking $\gamma_{b}/2\pi=10^{2}$ Hz, we obtain the entanglement between the cavity output mode and the magnon mode $E_{\mathrm{am}}=0.23$, which leads to the teleportation state fidelity $\mathcal{F}=0.55$, by choosing the central frequency $\Omega=-\omega_{b}$ (with respect to $\omega_{0}$, cf. Fig. 1(b)) and an optimal bandwidth $\tau^{-1}\approx 0.1\omega_{b}$ Genes2008 of the filter. Clearly, by properly filtering the output field, one realizes an effective entanglement distillation because the selected output mode is more entangled than the cavity mode with the magnon mode. The state fidelity $\mathcal{F}=0.55$ is above the minimum threshold of 0.5 for quantum teleportation with a coherent input state Braunstein2000 . We remark that, although the state fidelity can be high, e.g., up to 0.83, for CV quantum teleportation in purely optical systems Yukawa2008 ; Pirandola2015 ; Hu2023 , the fidelity in the CV quantum teleportation between objects of a different nature (e.g., light and matter) is typically much lower: the fidelity in the range of $0.58{-}0.60$ was demonstrated in the teleportation from light to an atomic ensemble Polzik2006 . Our fidelity $\mathcal{F}=0.55$ in the microwave-to-magnon quantum teleportation is comparable to the above values.
@@
Besides the applications in the channel multiplexing quantum teleportation, the work may be exploited for designing controllable quantum devices. In the chiral CMM system as illustrated in Fig. 1(a), the transmission direction of the circulating microwave cavity mode entangled with the magnon (phonon) mode can be conveniently switched on demand by reversing the bias magnetic field to change the precession direction of the magnetization Xie2023 . Such fast switch entanglement may find applications in relevant quantum protocols.
VI Conclusion

### v4

@@ 3 Nonreciprocal entanglements in the ideal case
We start with the case where the CW mode is driven. The CCW mode is essentially decoupled from the tripartite system of the CW, magnon, and mechanical modes, because of $g_{\circlearrowright}>0$ and $g_{\circlearrowleft}=0$. The system is then reduced to the tripartite CMM system and we adopt a red-detuned microwave drive field with respect to the magnon mode with detuning $\tilde{\Delta}_{m}\approx\omega_{b}$ (Fig. 1(b)), which is responsible for significantly cooling the low-frequency mechanical mode to eliminate the dominant thermal noise entering the system. The cooling interaction optimally works in the resolved-sideband limit of $\kappa_{m}\ll\omega_{b}$, which is well satisfied for the CMM experiments using YIG [8, 10, 11, 19]. A precondition for seeing entanglement in the system is that the mechanical mode should be cooled into its quantum ground state [9, 24, 25]. However, with only the effective magnomechanical cooling (beam-splitter) interaction, it is insufficient to produce entanglement. The magnomechanical parametric down conversion (PDC) interaction should be activated. Typically, this is realized by driving the magnon mode with a blue-detuned microwave field, which, however, can easily cause the system to be unstable. We thus adopt a sufficiently strong red-detuned drive field [87], such that the weak-coupling condition $G_{m}\ll\omega_{b}$ for taking the rotating-wave (RW) approximation to obtain the magnomechanical beam-splitter interaction $\propto m^{\dagger}b+mb^{\dagger}$ is broken, where $b$ denotes the mechanical mode and $b=(q+ip)/\!\sqrt{2}$. Consequently, the counter-RW terms, which correspond to the PDC interaction $\propto m^{\dagger}b^{\dagger}+mb$, start to play the role in creating magnomechanical entanglement. The entanglement is further distributed to the circulating cavity mode when the cavity is resonant with the Stokes mechanical sideband, i.e., $\Delta_{a}\approx-\omega_{b}$, yielding the entanglement between the CW microwave mode with the magnon (mechanical) mode, as shown in Fig. 2(a) (Fig. 2(b)). The entanglements are maximized around the optimal detunings as depicted in Fig. 1(b), i.e., the CW cavity and magnon modes are respectively resonant with the Stokes and anti-Stokes sidebands of the drive field.
@@
We have employed feasible parameters [8, 9, 10, 11, 19, 52]: $\omega_{a}/2\pi\approx\omega_{m}/2\pi\approx\omega_{0}/2\pi=10$ GHz, $\omega_{b}/2\pi=10$ MHz, $\kappa_{m}/2\pi=1$ MHz, $\gamma_{b}/2\pi=10^{2}$ Hz, $\kappa_{a,i}/2\pi=0.2$ MHz, $\kappa_{a,e}/2\pi=2.8$ MHz ($4.8$ MHz), $g_{\circlearrowright}/2\pi=4$ MHz ($8$ MHz), and $T=10$ mK in getting Fig. 2(a) (Fig. 2(b)). These parameters lead to a complex steady-state average of $\left\langle m\right\rangle$, cf. Eq. (5). Therefore, Fig. 2(a) (Fig. 2(b)) is plotted with a fixed amplitude of the effective coupling $|G_{m}|/2\pi=4.0$ MHz ($2.5$ MHz). Figure 2(c) shows the drive power required to generate considerable microwave-phonon entanglement for different values of the cavity decay rate. Clearly, one can increase the drive power to compensate the cavity losses in such a torus-shaped microwave cavity. Figure 2(d) plots the entanglements $E_{a_{\circlearrowright}m}$ and $E_{a_{\circlearrowright}b}$ versus the mechanical damping rate $\gamma_{b}$. Both the entanglements can survive even for a very large damping rate $\gamma_{b}/2\pi\sim 10^{4}$ Hz, corresponding to the mechanical quality factor $Q_{b}\sim 10^{3}$, which can be easily achieved in the CMM experiments [18]. Note that the results shown in Fig. 2 are in the steady state, which is guaranteed by the negative eigenvalues (real parts) of the drift matrix $A$. Under the parameters of Fig. 2(c) and taking $\kappa_{a}$, e.g. as used in Fig. 2(b), the maximum coupling strength $|G_{m}|$ for keeping the system stable is $2\pi\times 11.9$ MHz, which is much stronger than the value we used, so the system remains stable. Besides, a strong drive might lead to magnon frequency combs [20], which we check below. With the parameters of Fig. 2(c) and $\kappa_{a}$ used in Fig. 2(b), by investigating the time evolution of each mode of the system (Appendix A), we determine that the critical threshold of the coupling strength $|G_{m}|$ for producing the magnon frequency combs is about $2\pi\times 8.5$ MHz, which corresponds to a microwave driving power of $0.9$ W. The drive power we used is far below this value, so the presence of magnon frequency combs can be excluded. Note that the threshold power looks high, because we take a feasible relatively large cavity decay rate.
Figure 3: (a) [(b)]Tripartite entanglement in terms of the minimum residual contangle $R_{a_{\circlearrowright}mb}$ versus $\Delta_{a}$, with the paramters as used in plotting of Fig. 2(a) [Fig. 2(b)] and $\tilde{\Delta}_{m}=0.76\omega_{b}$ ($0.65\omega_{b}$).
Apart from bipartite entanglements studied above, we confirm that the genuine photon-magnon-phonon tripartite entanglement is also present in our system. The genuine tripartite entanglement is quantified by the minimum residual contangle $R_{\tau}^{\mathrm{min}}$ [89], which can be conveniently computed when the CM is achieved [9]. Figs. 3(a) and 3(b) respectively show the tripartite entanglement measure $R_{a_{\circlearrowright}mb}$ versus the detuning $\Delta_{a}$, under the parameters as used in Figs. 2(a) and 2(b).
@@
For the case where the CCW cavity mode is driven, due to the absence of the CCW mode-magnon coupling, i.e., $g_{\circlearrowleft}=0$, the magnon mode is effectively not driven, and hence the dispersive magnomechanical coupling is too weak to create any entanglement [18], i.e., $E_{a_{\circlearrowleft}m}=E_{a_{\circlearrowleft}b}=0$ and $R_{a_{\circlearrowleft}mb}=0$. Therefore, significant nonreciprocal microwave-magnon, microwave-phonon and microwave-magnon-phonon entanglements are achieved by exploiting the chiral cavity-magnon coupling in the system.
@@
4 Nonreciprocal entanglements versus experimental imperfections
In an actual experiment, various imperfections may cause appreciable effects on the nonreciprocity of the entanglements achieved in the ideal case (Sec. 3). These include the coupling between the two degenerate counter-propagating cavity modes by backscattering from the YIG sphere or the cavity ports [52], which can be modeled as $\hbar J(a_{\circlearrowright}a_{\circlearrowleft}^{\dagger}+a_{\circlearrowright}^{\dagger}a_{\circlearrowleft})$ with $J$ being the coupling strength, and the residual coupling of the Kittel mode to the CCW circulating mode, i.e., $g_{\circlearrowleft}\ll g_{\circlearrowright}$ but $g_{\circlearrowleft}\neq 0$, due to, e.g., the inaccurate position of the YIG sphere in the cavity. The inclusion of these two effects leads to a new Hamiltonian of the system and linearized QLEs for the quantum fluctuations, which are provided in the Appendix B for brevity.
Figure 4 manifests the impacts of the above two imperfections on the nonreciprocity of the entanglements. Specifically, in Fig. 4(a) we show the nonreciprocal microwave-magnon entanglement as a function of the mode coupling strength $J$ under $\chi\equiv g_{\circlearrowleft}/g_{\circlearrowright}=0.1$ ($g_{\circlearrowright}$ is fixed as in Fig. 2(a)). The four curves correspond, respectively, to the entanglement between the magnon mode with the CW or the CCW mode when either the CW or the CCW mode is driven. For the two solid curves with the CW mode under drive, the increase of the mode coupling $J$ leads to the distribution of the entanglement from the CW mode to the CCW mode, because they are coupled via the beam-splitter (state-swap) interaction, as witnessed by the increasing $E_{a_{\circlearrowleft}m}^{(\circlearrowright)}$ (the declining $E_{a_{\circlearrowright}m}^{(\circlearrowright)}$ ) with the coupling $J$ in Fig. 4(a). By contrast, for the two dashed curves with the CCW mode under drive, the CCW mode-magnon coupling $g_{\circlearrowleft}=0.1g_{\circlearrowright}$ is too weak (under the same power) to efficiently drive the magnon mode to obtain a sufficiently strong coupling $G_{m}$ for creating entanglement. Although the mode coupling also leads to the driving of the CW mode, this driving is typically weak and also inefficient to achieve a strong $G_{m}$ for generating considerable entanglement. Consequently, the entanglements $E_{a_{\circlearrowleft}m}^{(\circlearrowleft)}$ and $E_{a_{\circlearrowright}m}^{(\circlearrowleft)}$ are either zero or negligibly small. Similar results are observed for the microwave-phonon entanglement as shown in Fig. 4(b). We therefore conclude that the nonreciprocity of both the microwave-magnon and -phonon entanglements is robust against the backscattering induced mode coupling, even when the chiral coupling condition $g_{\circlearrowright}\gg g_{\circlearrowleft}=0$ is slightly broken.
@@ 4 Nonreciprocal entanglements versus experimental imperfections
Figure 5: Stationary entanglement $E_{a_{\circlearrowright}\nu}^{(\circlearrowright)}$ (thick solid), $E_{a_{\circlearrowleft}\nu}^{(\circlearrowright)}$ (thin solid), $E_{a_{\circlearrowright}\nu}^{(\circlearrowleft)}$ (thick dashed) and $E_{a_{\circlearrowleft}\nu}^{(\circlearrowleft)}$ (thin dashed) versus bath temperature $T$. Here, $\nu=m$ in (a) for microwave-magnon entanglement; $\nu=b$ in (b) for microwave-phonon entanglement. We take $\chi=0.1$ and the other parameters of (a) and (b) are the same as in Figs. 4(c) and 4(d), respectively.
In Fig. 5, we study the effect of thermal noises on the nonreciprocal entanglements. In general, the lower the bath temperature is, the stronger the entanglements are. Nevertheless, too low temperature also leads to the emergence of the other undesired entanglements. One thus can raise a bit the temperature, e.g., around 80 mK, to completely kill the unwanted entanglements, at the price of slightly reducing the desired entanglement. The nonreciprocal cavity-magnon and -phonon entanglements can be achieved for the temperature up to 125 mK and 163 mK, respectively, as shown in Figs. 5(a) and 5(b).
@@
In Fig. 6, we study the nonreciprocal tripartite entanglements versus the mode coupling $J$ (Fig. 6(a)) and the coupling ratio $\chi$ (Fig. 6(b)). The four curves represent the tripartite entanglement of the magnon mode, the mechanical mode, and the CW or the CCW mode for the two situations when the CW and CCW modes are respectively driven. The nonreciprocity of the tripartite entanglement is clearly shown to be robust against the mode coupling $J$ and the residual coupling $g_{\circlearrowleft}$ to the CCW mode, as witnessed by the vanishing residual contangle $R_{\tau}^{\mathrm{min}}\approx 0$ of the three unwanted tripartite entanglements.
5 Potential applications
The proposed nonreciprocal entanglements find many potential applications, one of which is to realize the channel multiplexing quantum teleportation, as depicted in Fig. 7. In the conventional continuous-variables (CV) quantum teleportation scheme (Fig. 7(a)) [90], the teleportation of a quantum state from the sender onto the quantum memory exploits a single channel. By contrast, the nonreciprocal entanglement allows for a channel multiplexing protocol (Fig. 7(b)), where the chiral cavity-magnon coupling enables the microwave photons injecting from the left (right) side to couple with the magnomechanical system in the right (left) cavity and entangle with the magnons/phonons contained therein, but decouple from the magnomechanical system in the left (right) cavity. This allows for the teleportation of the quantum state from the senders at both sides onto their respective quantum memories still using a single channel, thereby increasing the usage efficiency of the channel compared to the conventional teleportation scheme.
Figure 7: (a) Schematic diagram of the conventional CV quantum teleportation. By exploiting the prepared Einstein-Podolsky-Rosen (EPR) state, the “Bell-state" measurement (realized via the homodyne detection) teleports the quantum state from the sender to the quantum memory [90]. (b) The channel multiplexing quantum teleportation protocol based on the nonreciprocal entanglement proposed in this work. See text for description.
We now discuss the performance of such a channel multiplexing protocol based on, e.g., the generated nonreciprocal microwave-magnon entanglement, and calculate the corresponding teleportation state fidelity. In the CV quantum teleportation, the fidelity of teleporting an input single-mode Gaussian state by exploiting a two-mode Gaussian entangled state is given by [91]
@@ 5 Potential applications
where $V=2V_{in}+\sigma_{z}V_{e}\sigma_{z}^{T}+\sigma_{z}V_{ef}+V_{ef}^{T}\sigma_{z}^{T}+V_{f}$, with $\sigma_{z}=\mathrm{diag}[1,-1]$ being the Pauli-$z$ matrix, $V_{in}$ being the CM of the input state, and $V_{e}$, $V_{f}$, $V_{ef}$ and $V_{fe}$ being $2\times 2$ blocks of the CM of the two-mode entangled state. To be specific, we consider a microwave coherent state as the input state with its CM $V_{in}=\mathrm{diag}[1/2,1/2]$, and the generated microwave-magnon entangled state as the two-mode entangled state. It should be noted that the above microwave-magnon entanglement is between the microwave cavity mode and the magnon mode, while the subsequent “Bell-state" measurement is applied onto the cavity output field and the input coherent state. To estimate the fidelity using Eq. (8), one has to properly define the cavity output mode to have a physically valid CM. The cavity output mode can be defined with a filter function with a properly chosen central frequency $\Omega$ and bandwidth $1/\tau$ (F1 and F2 in Fig. 7(b) denote two filters), given by [92]
@@
Under the parameters of Fig. 2(d) and taking $\gamma_{b}/2\pi=10^{2}$ Hz, we obtain the entanglement between the cavity output mode and the magnon mode $E_{\mathrm{am}}=0.23$, which leads to the teleportation state fidelity $\mathcal{F}=0.55$, by choosing the central frequency $\Omega=-\omega_{b}$ (with respect to $\omega_{0}$, cf. Fig. 1(b)) and an optimal bandwidth $\tau^{-1}\approx 0.1\omega_{b}$ [92] of the filter. The state fidelity $\mathcal{F}=0.55$ is above the minimum threshold of 0.5 for quantum teleportation with a coherent input state [93]. We remark that, although the state fidelity can be high, e.g., up to 0.83, for CV quantum teleportation in purely optical systems [94, 95, 96], the fidelity in the CV quantum teleportation between objects of a different nature (e.g., light and matter) is typically much lower: the fidelity in the range of $0.58{-}0.60$ was demonstrated in the teleportation from light to an atomic ensemble [97]. Our fidelity $\mathcal{F}=0.55$ in the microwave-to-magnon quantum teleportation is comparable to the above values.
@@
Besides the applications in the channel multiplexing quantum teleportation, the work may be exploited for designing controllable quantum devices. In the chiral CMM system as illustrated in Fig. 1(a), the transmission direction of the circulating microwave cavity mode entangled with the magnon (phonon) mode can be conveniently switched on demand by reversing the bias magnetic field to change the precession direction of the magnetization [71]. Such fast switch entanglement may find applications in relevant quantum protocols.
6 Conclusion